OpenStudy (anonymous):

I dont understand this question: Rewrite each of the following in the form of a (x+p)^2+q where a, p and q belongs to real numbers: 1) 3x^2+6x-4 2)-2x^2-8x+4

OpenStudy (anonymous):

this question is asking u to use complete square method

OpenStudy (anonymous):

i dont know how to do that method

OpenStudy (anonymous):

\[ax^2+bx+c=a(x-p)^2+q\] where \[p =-\frac{b}{2a} \space and \space q = c-\frac{b^2}{4ac}\]

OpenStudy (anonymous):

so look at ur first equation 3x^2+6x-4, this means a=3, b=6 and c =-4

OpenStudy (anonymous):

put those numbers in the above formulas u'll get the required answer

OpenStudy (anonymous):

p=-6/(2.3)=-1, and q=-4-((36/4.3.(-4))

OpenStudy (anonymous):

this is one way to do it, the other way to solve this problem is by doing the following

OpenStudy (anonymous):

3x^2+6x-4=0, so divide both sides by the coefficient of x^2 which is 3

OpenStudy (anonymous):

x^2+2x-4/3=0

OpenStudy (anonymous):

so are the p and q values -1 and -112?

OpenStudy (anonymous):

well if u did as i told u then yes

OpenStudy (anonymous):

so if those are the values then i just rearrange the question into the form of a(x+p)^2+q?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

so for the first question would it be, 3(x+-1)^2+-112?

OpenStudy (anonymous):

you mean -112

OpenStudy (anonymous):

yes no plus sign

OpenStudy (anonymous):

3(x-1)^2-112?

OpenStudy (anonymous):

yeah thats it right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so try doing question 2

OpenStudy (anonymous):

i do exactly the same thing for question 2 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is the p value for question 2 = -2?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and is the q value = 2?

OpenStudy (anonymous):

because thats what i calculated